Composite scores on the ACT have an average of 20.6 and a variance of 36.
Calculate the z-score for a student with a composite ACT score of 25.1. Round your answer to two decimal places.

Answer by math_tutor2020(3835)


mu = 20.6 = mean
sigma = 6 = standard deviation

note:
sigma = sqrt(variance) = sqrt(36) = 6

Let's compute the z score for x = 25.1
z = (x - mu)/sigma
z = (25.1 - 20.6)/6
z = 4.5/6
z = 0.75

Answer: 0.75